OpenStudy (anonymous):

h(x)=x^2+3x/4x+27 find h(-8) can someone please help?!?!?!

OpenStudy (anonymous):

so what the problem is telling you x=-8 so fill in your variables and solve

OpenStudy (anonymous):

thank you.. would it be -8= -8^2+3(8)/4(8)+27?

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

to me it looks like a false answer, is that an option?

OpenStudy (anonymous):

it says to simplify my answer is all..

OpenStudy (anonymous):

well nvm lol just simplify

OpenStudy (anonymous):

this is harder than I thought it'd be. I got -6.625 but it says its incorrect

OpenStudy (anonymous):

\(f(-8)\) does NOT mean set it equal to \(-8\) and solve for \(x\) it means replace \(x\) by \(-8\) and see what you get

OpenStudy (anonymous):

\[h(x)=\frac{x^2+3x}{4x+27}\]\[h(-8)=\frac{(-8)^2+3\times (-8)}{4\times (-8)+27}\]